These two waves travel along the same string:

y1 = (4.10 mm) sin(1.51?x - 370?t)
y2 = (5.53 mm) sin(1.51?x - 370?t + 0.867?rad).

What are (a) the amplitude and (b) the phase angle (relative to wave 1) of the resultant wave? (c) If a third wave of amplitude 5.11 mm is also to be sent along the string in the same direction as the first two waves, what should be its phase angle in order to maximize the amplitude of the new resultant wave?
Answer #1

The equation for wave first is, a sin(tta) And the second wave is, 2ah sin (a) The resultant wave is the sum of the two waves

The resultant amplitude of the wave is, a -Va^+a,+2aa, cos(a-) 4.10 mm) +(5.53 mm)+2(4.10 mm)(5.53 mm)cos(0-0.867T) (4.10 mm) +(5.53 m)-41.419 2.443 mm

The resultant phase angle is, (4.10 mm )sin( 0)4(5.53 mm )sin( 0.867r) (4.10 mm)cos(0)+(5.53 mm)cos(0.867/7) tanの 2.250 4.10-5.051 ga = tan-1 (-2.3659) 1.170 = 0.37287 The phase shift related to the first wave is, phase shift = 0.3728-0 = 0.37287

The addition of the third wave is the superposition of the resultant wave and the third wave. If the third wave has the same frequency and wave number as that of the first two waves. Therefore, Thus, the resultant amplitude of the resultant wave and the third wave is, The phase of the resultant wave is effected by cosine function only To maximize the phase angle, the cosine function should be equal to 1 Therefore, cos(2-)-1 0.3728T